1 three numbers are in the ratio 3 4 5 if the sum of the largest and the smallest equals the sum of the third and 52, find the number

2 the difference between two integer 9 and the ratio of these integers is 4:5. find these integers. solution

Answer by ewatrrr(24785)


Hi
Let x represent the constant of proportionality
1) 3x + 5x = 4x + 52
4x = 52
x = 13
Numbers: 39, 52, 65
And...
39 + 65 = 52 + 52 CHECKS!